# Tackling Negative Transfer on Graphs

### v1

@@ Abstract
Transfer learning aims to boost the learning on the target task leveraging knowledge learned from other relevant tasks. However, when the source and target are not closely related, the learning performance may be adversely affected, a phenomenon known as negative transfer. In this paper, we investigate the negative transfer in graph transfer learning, which is important yet underexplored. We reveal that, unlike image or text, negative transfer commonly occurs in graph-structured data, even when source and target graphs share semantic similarities. Specifically, we identify that structural differences significantly amplify the dissimilarities in the node embeddings across graphs. To mitigate this, we bring a new insight: for semantically similar graphs, although structural differences lead to significant distribution shift in node embeddings, their impact on subgraph embeddings could be marginal. Building on this insight, we introduce two effective yet elegant methods, Subgraph Pooling (SP) and Subgraph Pooling++ (SP++), that transfer subgraph-level knowledge across graphs. We theoretically analyze the role of SP in reducing graph discrepancy and conduct extensive experiments to evaluate its superiority under various settings. Our code and datasets are available at: https://github.com/Zehong-Wang/Subgraph-Pooling.
@@ 1 Introduction
Graph Neural Networks (GNNs) are widely employed for graph mining tasks across various fields Gaudelet et al. (2021); Kipf and Welling (2017); He et al. (2020). Despite their remarkable success in graph-structured datasets, these methods exhibit limitations in label sparse scenarios Dai et al. (2022a). This restricts the applications of GNNs in real-world datasets where label acquisition is challenging or impractical. To address the issue, transfer learning Zhuang et al. (2020) emerges as a solution, which aims to transfer knowledge from a label-rich source graph to a label-sparse target graph through fine-tuning or prompting Sun et al. (2023).
However, the success of transfer learning is not always guaranteed Wang et al. (2019); Zhang et al. (2022). If the source and target lack sufficient similarity, transferring knowledge from such weakly related source may impair performance on the target. This phenomenon is known as negative transfer, which was initially analyzed by Wang et al. (2019). By interpreting transfer learning as a specific generalization problem, it demonstrated that negative transfer is caused by the divergence between joint distributions of the source and target. To this end, researchers employed adversarial learning Wu et al. (2020), causal learning Chen et al. (2022), or domain regularizer You et al. (2023) to develop domain-invariant encoders, which reduce the distribution shift between the source and target.
@@
In this paper, we conduct a systematic analysis of the negative transfer issue in GNNs, which is a lack of existing works. Our observations indicate that in graph datasets, negative transfer often occurs even when the source and target are semantically similar. This is in contrast to image and text datasets, where similar sources typically enhance the performance on targets Zhuang et al. (2020). We consider this derives from the differences in graph structures between the source and target, which may lead to significant distribution shifts on node embeddings. For example, in financial transaction networks collected over different time intervals, transaction patterns can vary markedly due to the impact of social events or policy changes. These evolving patterns notably change the local structure of users, leading to a substantial divergence in user embeddings. To address this challenge, we introduce two straightforward yet effective methods called Subgraph Pooling (SP) and Subgraph Pooling++ (SP++) to reduce the discrepancy between graphs. Our major contributions are summarized as follows:
  • Negative Transfer in GNNs. We systematically analyze the negative transfer in GNNs. We find that the structural difference between the source and target graphs intensifies distribution shifts on node embeddings, as the aggregation process of GNNs is highly sensitive to perturbations in graph structures. To address this issue, we present a novel insight: for semantically similar graphs, although structural differences lead to significant distribution shift in node embeddings, their impact on subgraph embeddings could be marginal.
  • Subgraph Pooling to Tackle Negative Transfer. Building upon this insight, we introduce plug-and-play modules Subgraph Pooling (SP) and Subgraph Pooling++ (SP++) to mitigate the negative transfer. The key idea is to transfer subgraph information across source and target to prevent the distribution shift. Notably, we provide a comprehensive theoretical analysis to clarify the operational processes behind Subgraph Pooling.
@@ 2.2 Why Negative Transfer Happens?
Although negative transfer typically occurs between two weakly related domains, it remains a prevalent issue in GNNs, even when the source and target are semantically similar. This issue is attributed to the sensitivity of GNNs to graph structures. Specifically, differences in structural distribution between the source and target can lead to distinct marginal distributions $P_{S}(\mathcal{Z})$ and $P_{T}(\mathcal{Z})$, thereby diverging the corresponding joint distributions. To support this claim, we analyze the influence of structures on graph discrepancy (CMD value) and transfer learning performance, as illustrated in Figure 1 (Left). Particularly, when graph structure is not considered (using MLP), the discrepancy remains relatively low, ensuring the performance gain of transfer learning. Conversely, incorporating structural information through GNNs can increase the discrepancy, resulting in negative transfer.
@@
2.3 Analysing The Impact of Structure
@@ 3.1 Subgraph Pooling
We start by presenting the final objective for node-level graph transfer learning. The goal is to minimize the empirical risk (loss) on the target distribution, expressed as
@@
where $P_{T}(\mathcal{Z},\mathcal{Y})$ represents the joint distribution over the target graph, $z$ denotes the node embeddings encoded by a GNN backbone $f_{\Theta}(\cdot)$ with parameters $\Theta$, and $g(\cdot)$ is a linear classifier. These two components are represented as
@@
However, owing to the scarcity of labels, we cannot describe the real joint distribution $P_{T}(\mathcal{Z},\mathcal{Y})$ Wenzel et al. (2022). Consequently, directly optimizing Eq. 4 on the target graph may lead to overfitting Mallinar et al. (2022). To address this, we pre-train the backbone and classifier on a similar, label-rich source graph to achieve a suitable initialization, aiming to project the source and target into similar embedding spaces. Nevertheless, training such encoder is challenging, as structural differences enlarge the disparity between the marginal distributions $P_{S}(\mathcal{Z})$ and $P_{T}(\mathcal{Z})$. To overcome the limitation, we introduce Subgraph Pooling (SP), a plug-and-play method that leverages subgraph information to diminish the discrepancy between the source and target. This approach is based on the following assumption.
@@
where $\mathbf{h}_{i}\in\mathbf{H}$ represents the subgraph embeddings (the new embeddings for each node), utilized in training the classifier $g(\cdot)$. $w_{ij}$ denotes the pooling weight, which can be either learnable or fixed. Empirically, the MEAN pooling function is effective enough to achieve desirable transfer performance.
The integration of the SP layer into GNN architectures does not substantially increase time complexity. We have three-fold considerations. Firstly, the SP layer functions as a non-parametric GNN layer, thus imposing no additional burden on model optimization and enjoying the computational efficiency of existing GNN libraries Fey and Lenssen (2019). Secondly, the sampling operation relies solely on the graph structure and can be performed in pre-processing. Finally, our empirical observations indicate that sampling low-order neighborhoods ($k=1,2$) is sufficient for achieving optimal transfer learning performance, which ensures computational efficiency in both sampling and pooling.
Our SP layer leverages subgraph information to reduce the discrepancy (CMD value) between node embeddings in the source and target, thereby enhancing transfer learning performance, which is illustrated in Figure 1 (Left). Furthermore, the use of subgraph information also reduces the sensitivity to structural perturbations, as evidenced in Figure 1 (Right). We also provide a theoretical analysis to explain how subgraph information reduces the graph discrepancy.
@@ 3.2 Subgraph Pooling++
The performance of Subgraph Pooling highly depends on the choice of subgraph sampling function. Employing an inappropriate sampling function can impair the distinguishability of the learned embeddings. For instance, the basic $k$-hop sampler might cause distinct nodes to share an identical subgraph, collapsing the embeddings into a single point. This can result in the potential over-smoothing problem Zhao and Akoglu (2020); Keriven (2022); Huang et al. (2023) since these nodes may have different labels.
To address the limitation, we propose an advanced method called Subgraph Pooling++ (SP++) that leverages Random Walk (RW) Huang et al. (2021) to sample subgraphs. We use the same hyper-parameter $k$ to define the maximum walk length, restricting the sampling process within $k$-hop subgraphs. The RW sampler is defined as
@@
Another approach to mitigate the risk of over-smoothing is to design advanced pooling functions. For example, we can employ attention mechanism Lee et al. (2019) or hierarchical pooling Wu et al. (2022a) to adaptively assign pooling weights $w_{ij}$ to nodes within a subgraph, thus preserving the uniqueness of subgraph embeddings. However, empirical evidence suggests that these advanced methods offer no significant advantage compared to basic MEAN pooling (Sec. 4.3). Moreover, there is a concern regarding the efficiency of complicated pooling functions, as they could potentially increase computational and optimization efforts at each epoch. In contrast, the proposed RW sampling can be efficiently executed during pre-processing. To illustrate how RW sampling alleviates over-smoothing, we provide a concrete example below.
@@
Illustration. Let $\mathbf{h}_{u}=\sum_{i\in\mathcal{N}_{s}(u)\cup u}\mathbf{z}_{i}/(|\mathcal{N}_{s}(u)|+1)$ and $\mathbf{h}_{v}=\sum_{j\in\mathcal{N}_{s}(v)\cup v}\mathbf{z}_{j}/(|\mathcal{N}_{s}(v)|+1)$ as subgraph embeddings for node $u,v$ via MEAN pooling, where $\mathbf{h}_{u}=\mathbf{h}_{v}$. The empirical risk with classifier $g(\cdot)$ is given by:
@@
For simplicity, we use the mean square loss. We consider that $R_{S}$ is minimized when $g(\mathbf{h}_{u})=y_{u}$ and $g(\mathbf{h}_{v})=y_{v}$, but it is impossible to find a classifier $g(\cdot)$ that projects a single vector into different labels. However, by applying RW to sample subgraphs with $\mathcal{N}_{r}(u)$ and $\mathcal{N}_{r}(v)$, we can obtain distinct subgraph embeddings $\overline{\mathbf{h}}_{u}\neq\overline{\mathbf{h}}_{v}$. Then, it becomes feasible to find a classifier $g^{*}(\cdot)$ such that $g^{*}(\overline{\mathbf{h}}_{u})=y_{u}$ and $g^{*}(\overline{\mathbf{h}}_{v})=y_{v}$ to minimize $R_{S}$. In the extreme case, the subgraphs sampled by RW might be the same as $k$-hop, leading to $\overline{\mathbf{h}}_{u}=\overline{\mathbf{h}}_{v}$. To mitigate the issue, we can control the walk length and sampling frequency to maintain the distinctiveness of the sampled subgraphs. Therefore, utilizing RW sampler can lead to a lower empirical risk.
@@ 4.1 Experimental Setup
Datasets. We use Citation network Wu et al. (2020), consisting of ACMv9 and DBLPv8; Airport network Ribeiro et al. (2017), including Brazil, USA, and Europe; Twitch network Rozemberczki et al. (2021) collected from six countries, including DE, EN, ES, FR, PT, RU; citation network Arxiv Hu et al. (2020) consisting papers with varying publish times; and dynamic financial network Elliptic Weber et al. (2019) that contains dozens of graph snapshots where each node is a Bitcoin transaction.
Baselines. We use four different GNNs as backbones: GCN Kipf and Welling (2017), SAGE Hamilton et al. (2017), GAT Veličković et al. (2018), and SGC Wu et al. (2019). We compare our proposed Subgraph Pooling with No Transfer (directly training on target), Empirical Risk Minimization (pre-training on source and fine-tuning on target), Multi-task (jointly training on source and target), EERM Wu et al. (2022b), and recent SOTA method GTrans Jin et al. (2023). We also compare various domain adaptation methods, including DANN Ganin et al. (2016), CDAN Long et al. (2018), UDAGCN Wu et al. (2020), MIXUP Wang et al. (2021), EGI Zhu et al. (2021b), SR-GNN Zhu et al. (2021a), GRADE Wu et al. (2023a), SSReg You et al. (2023), and StruRW Liu et al. (2023a).
Settings. We pre-train the model on the source with 60 percent of labeled nodes and adapt the model to the target. The adaptation involves three settings: (1) directly applying the pre-trained model without any fine-tuning (Without FT); (2) fine-tuning the last layer (classifier) of the model (FT Last Layer); (3) fine-tuning all parameters (Fully FT). We take split 10/10/80 to form train/valid/test sets on the target graph. Specific hyper-parameters are presented in the Appendix C.
@@ 4.2 Node Classification
One-to-One Transfer. We use Citation and Airport as benchmarks and employ 2-layer GNNs with 64 dimensions as backbones. Considering the limited number of parameters, we apply the FT Last Layer setting. The transfer learning results on Citation over three GNN backbones are presented in Table 2. Our proposed GNN-SP outperforms other baselines across all settings and is even better than the model trained from scratch (No Transfer), demonstrating the capability of overcoming negative transfer. Note that GNN-SP performs better than the advanced GNN-SP++ with extremely limited labels. It is may because (1) the dataset is small and sparse ($\sim$5k to $\sim$7k nodes and $\sim$20k to $\sim$30k edges) where RW sampler fails to model the real localized structures; and (2) the label sparsity exacerbates overfitting.
The performance on Airport is presented in Table 3. Our method surpasses all baselines and achieves an average Rank of 1.5, which is notably higher than the No Transfer. Note that transferring knowledge to USA results in negative transfer for all baselines, likely due to its significantly larger size compared to the other two graphs. Specifically, USA contains 1,190 nodes and 28,388 edges, whereas Europe and Brazil contain only 399 nodes & 12,385 edges, and 131 nodes & 2,137 edges, respectively. The smaller graphs provide limited patterns and may introduce unexpected biases during pre-training. Additionally, we observe that GNN-SP++ performs better than GNN-SP, likely because these graphs are densely connected, leading to the issue that various nodes share identical $k$-hop subgraphs.
Figure 4: Node classification performance on Elliptic.
@@
One-to-Multi Transfer. We use Twitch as the benchmark, which consists of six graphs with different sizes and data distributions. We pre-train the model on DE and fine-tune on other graphs (EN, ES, FR, PT, RU). We employ ROC-AUC as metric and adopt 2-layer GCN and SAGE as backbones. Figure 3 shows GNN-SP outperforms standard GNN with up to 8% improvements on ROC-AUC under FT Last Layer setting and achieves better performance than the model directly trained on the target over 10 out of 12 settings. The results validate the generalizability of GNN-SP to multiple graphs.
Transfer with Dynamic Shift. In this scenario, the model is pre-trained on datasets collected from the past, and then fine-tuned on future data to evaluate its capability in handling temporal distribution shifts. We first adopt a dynamic financial network Elliptic with splitting 5/5/33 snapshots for train/valid/test and F1-score for evaluation. Figure 4 presents the results where the test snapshots are grouped into 9 folds in chronological order. Our GNN-SP outperforms standard GNN up to 10% and 24% improvements over GAT and SAGE backbones, respectively.
@@
Additionally, we use Arxiv as another temporal dataset, where nodes represent papers published from 2005 to 2020 and edges indicate citations. Based on the publication time, we collect five sub-graphs, represented as Time 1 (2005 - 2007), Time 2 (2008 - 2010), Time 3 (2011 - 2014), Time 4 (2015 - 2017), and Time 5 (2018 - 2020). We use the first four graphs as sources and the last one as the target. The results are presented in Table 4 where our GNN-SP++ achieves significant improvements over all baselines. We note that the temporal distribution shift is marginal when the source and target are temporally proximate, resulting in improved transfer learning performance. Additionally, the performance of GNN-SP++ is considerably superior to GNN-SP, further validating the efficacy of the RW sampler.

### v2

# Subgraph Pooling: Tackling Negative Transfer on Graphs

@@ Abstract
Transfer learning aims to enhance performance on a target task by using knowledge from related tasks. However, when the source and target tasks are not closely aligned, it can lead to reduced performance, known as negative transfer. Unlike in image or text data, we find that negative transfer could commonly occur in graph-structured data, even when source and target graphs have semantic similarities. Specifically, we identify that structural differences significantly amplify the dissimilarities in the node embeddings across graphs. To mitigate this, we bring a new insight in this paper: for semantically similar graphs, although structural differences lead to significant distribution shift in node embeddings, their impact on subgraph embeddings could be marginal. Building on this insight, we introduce Subgraph Pooling (SP) by aggregating nodes sampled from a k-hop neighborhood and Subgraph Pooling++ (SP++) by a random walk, to mitigate the impact of graph structural differences on knowledge transfer. We theoretically analyze the role of SP in reducing graph discrepancy and conduct extensive experiments to evaluate its superiority under various settings. The proposed SP methods are effective yet elegant, which can be easily applied on top of any backbone Graph Neural Networks (GNNs). Our code and data are available at: https://github.com/Zehong-Wang/Subgraph-Pooling.
@@ 1 Introduction
Graph Neural Networks (GNNs) are widely employed for graph mining tasks across various fields Gaudelet et al. (2021); Kipf and Welling (2017); He et al. (2020). Despite their remarkable success in graph-structured datasets, these methods exhibit limitations in label sparse scenarios Dai et al. (2022a). This restricts the applications of GNNs in real-world datasets where label acquisition is challenging or impractical. To address the issue, transfer learning Zhuang et al. (2020) emerges as a solution, which aims to transfer knowledge from a label-rich source graph to a label-sparse target graph through fine-tuning or prompting.
However, the success of transfer learning is not always guaranteed Wang et al. (2019); Zhang et al. (2022). If the source and target tasks are not closely aligned, transferring knowledge from such weakly related sources may impair the performance on the target, known as negative transfer Wang et al. (2019). By interpreting transfer learning as a generalization problem, Wang et al. (2019) demonstrated that negative transfer derives from the divergence between joint distributions of the source and target tasks. To this end, researchers employed adversarial learning Wu et al. (2020), causal learning Chen et al. (2022), or domain regularizer You et al. (2023) to develop domain-invariant encoders, reducing the distribution shift between the source and target.
@@
In this paper, we systematically analyze the negative transfer on graphs, a lack of existing works. We find negative transfer often occurs in graph datasets even when the source and target graphs are semantically similar. The observation contrasts with visual and textual modalities, where similar sources typically enhance the performance on targets Zhuang et al. (2020). We identify that the issue stems from the graph structural differences between the source and target graphs, which leads to significant distribution shifts on node embeddings. For example, in financial transaction networks Weber et al. (2019) collected over different time intervals, the patterns of transactions can vary significantly due to the influence of social events or policy changes. These evolving patterns notably alter the local structure of users, leading to divergence in user embedding. To tackle the negative transfer on graphs, we introduce two straightforward yet effective methods called Subgraph Pooling (SP) and Subgraph Pooling++ (SP++) to reduce the discrepancy between graphs. Our major contributions are summarized as follows:
  • Negative Transfer in GNNs. We systematically analyze the negative transfer on graphs. We find that the structural difference between the source and target graphs amplifies distribution shifts on node embeddings, as the aggregation process of GNNs is highly sensitive to perturbations in graph structures. To address this issue, we present a novel insight: for semantically similar graphs, although structural differences lead to significant distribution shift in node embeddings, their impact on subgraph embeddings could be marginal.
  • Subgraph Pooling to Tackle Negative Transfer. Building upon this insight, we introduce plug-and-play modules Subgraph Pooling (SP) and Subgraph Pooling++ (SP++) to mitigate the negative transfer. The key idea is to transfer subgraph information across source and target graphs to prevent the distribution shift. Notably, we provide a comprehensive theoretical analysis to clarify the mechanisms behind Subgraph Pooling.
@@ 2.2 Why Negative Transfer Happens?
Negative transfer occurs in GNNs even if the source and target graphs are semantically similar. This issue is attributed to the sensitivity of GNNs on graph structures, where different structures diverge marginal distributions $P_{S}(\mathcal{Z})$ and $P_{T}(\mathcal{Z})$ between the source and target graphs. To support this claim, we illustrate the impact of structural differences on distribution discrepancy (CMD value) and transfer learning performance in Figure 1 (Left). Particularly, the discrepancy remains low if graph structure is ignored (using MLP), ensuring the performance gain of transfer learning. Conversely, incorporating structural information through GNNs can increase the discrepancy, resulting in negative transfer.
@@
2.3 Analyzing The Impact of Structure
@@ 3.1 Subgraph Pooling
The goal of node-level graph transfer learning is to minimize the empirical risk (loss) on the target distribution:
@@
where $P_{T}(\mathcal{Z},\mathcal{Y})$ is the joint distribution over the target graph, $z$ denotes the node embeddings encoded by a GNN backbone $f_{\Theta}(\cdot)$ with parameters $\Theta$, and $g(\cdot)$ is a linear classifier:
@@
However, owing to the scarcity of labels, we cannot exactly describe the joint distribution $P_{T}(\mathcal{Z},\mathcal{Y})$ Wenzel et al. (2022). Consequently, directly optimizing Eq. 4 on the target graph may lead to overfitting Mallinar et al. (2022). Additionally, based on the above discussion, it is non-trivial to obtain a suitable initialization by pre-training the model on a semantically similar, label-rich source graph, as structural differences enlarge the discrepancy between the marginal distributions $P_{S}(\mathcal{Z})$ and $P_{T}(\mathcal{Z})$. To overcome the limitation, we introduce Subgraph Pooling (SP), a plug-and-play method that leverages subgraph information to diminish the discrepancy between the source and target. This approach is based on the following assumption.
@@
where $\mathbf{h}_{i}\in\mathbf{H}$ represents the subgraph embeddings (the new embeddings for each node), utilized in training the classifier $g(\cdot)$. $w_{ij}$ denotes the pooling weight, which can be either learnable or fixed. Empirically, the MEAN pooling function is effective enough to achieve desirable transfer performance. By leveraging subgraph information, the SP layer reduces the discrepancy (CMD value) between node embeddings in the source and target, thereby enhancing transfer learning performance, illustrated in Figure 1 (Left). Furthermore, this also reduces the sensitivity on structural perturbations, as evidenced in Figure 1 (Right).
The integration of the SP layer into GNN architectures does not substantially increase time complexity. We have three-fold considerations. Firstly, the SP layer functions as a non-parametric GNN layer, thus imposing no additional burden on model optimization and enjoying the computational efficiency of existing GNN libraries Fey and Lenssen (2019). Secondly, the sampling operation relies solely on the graph structure and can be performed in pre-processing. Finally, our empirical observations indicate that sampling low-order neighborhoods ($k=1,2$) is sufficient for achieving optimal transfer learning performance, which ensures computational efficiency in both sampling and pooling. We also provide a theoretical analysis to explain how subgraph information reduces the graph discrepancy.
@@ 3.2 Subgraph Pooling++
The performance of Subgraph Pooling highly depends on the choice of subgraph sampling function. Employing an inappropriate sampling function can impair the distinguishability of the learned embeddings. For instance, the basic $k$-hop sampler might cause distinct nodes to share an identical subgraph, collapsing the embeddings into a single point. This can result in the potential over-smoothing Zhao and Akoglu (2020); Keriven (2022); Huang et al. (2023).
To address the limitation, we propose an advanced method Subgraph Pooling++ (SP++) that leverages Random Walk (RW) Huang et al. (2021) to sample subgraphs. We use the same hyper-parameter $k$ to define the maximum walk length, restricting the sampling process within $k$-hop subgraphs. The RW sampler is defined as
@@
Another approach to mitigate the risk of over-smoothing is to design advanced pooling functions. For example, we can employ attention mechanism Lee et al. (2019) or hierarchical pooling Wu et al. (2022a) to adaptively assign pooling weights $w_{ij}$ to nodes within a subgraph, thus preserving the uniqueness of subgraph embeddings. However, empirical evidence suggests that these pooling methods cannot offer significant advantages over basic MEAN pooling (Sec. 4.3). Moreover, there is a concern regarding the efficiency of complicated pooling functions, as they could potentially increase computational and optimization efforts at each epoch. In contrast, the proposed RW sampling can be efficiently executed during pre-processing. To illustrate how RW sampling alleviates over-smoothing, we provide a concrete example below.
@@
Illustration.
Let $\mathbf{h}_{u}=\sum_{i\in\mathcal{N}_{s}(u)\cup u}\mathbf{z}_{i}/(|\mathcal{N}_{s}(u)|+1)$ and $\mathbf{h}_{v}=\sum_{j\in\mathcal{N}_{s}(v)\cup v}\mathbf{z}_{j}/(|\mathcal{N}_{s}(v)|+1)$ as subgraph embeddings for node $u,v$ via MEAN pooling, where $\mathbf{h}_{u}=\mathbf{h}_{v}$. The empirical risk with classifier $g(\cdot)$ is given by:
@@
For simplicity, we use the mean square loss. $R_{S}$ is minimized when $g(\mathbf{h}_{u})=y_{u}$ and $g(\mathbf{h}_{v})=y_{v}$, but it is impossible to find a classifier $g(\cdot)$ to project a single vector into different labels. However, by applying RW to sample subgraphs with $\mathcal{N}_{r}(u)$ and $\mathcal{N}_{r}(v)$, we can obtain distinct subgraph embeddings $\overline{\mathbf{h}}_{u}\neq\overline{\mathbf{h}}_{v}$. Then, it becomes feasible to find a classifier $g^{*}(\cdot)$ satisfying $g^{*}(\overline{\mathbf{h}}_{u})=y_{u}$ and $g^{*}(\overline{\mathbf{h}}_{v})=y_{v}$. In the extreme case, the subgraphs sampled by RW might be the same as $k$-hop, leading to $\overline{\mathbf{h}}_{u}=\overline{\mathbf{h}}_{v}$. To mitigate the issue, we can control the walk length and sampling frequency to maintain the distinctiveness of the sampled subgraphs. Therefore, utilizing RW sampler can lead to a lower empirical risk.
@@ 4.1 Experimental Setup
Datasets.
We use Citation Wu et al. (2020), consisting of ACMv9 and DBLPv8; Airport Ribeiro et al. (2017), including Brazil, USA, and Europe; Twitch Rozemberczki et al. (2021) collected from six countries, including DE, EN, ES, FR, PT, RU; Arxiv Hu et al. (2020) consisting papers with varying publish times; and dynamic financial network Elliptic Weber et al. (2019) that contains dozens of graph snapshots where each node is a Bitcoin transaction.
Baselines.
We include four GNN backbones: GCN Kipf and Welling (2017), SAGE Hamilton et al. (2017), GAT Veličković et al. (2018), and SGC Wu et al. (2019). We compare our SP with No Transfer (directly training on target), Empirical Risk Minimization (transferring knowledge from source to target), Multi-task (jointly training on source and target), EERM Wu et al. (2022b), and recent SOTA method GTrans Jin et al. (2023). We also compare various domain adaptation methods, including DANN Ganin et al. (2016), CDAN Long et al. (2018), UDAGCN Wu et al. (2020), MIXUP Wang et al. (2021), EGI Zhu et al. (2021b), SR-GNN Zhu et al. (2021a), GRADE Wu et al. (2023a), SSReg You et al. (2023), and StruRW Liu et al. (2023a).
Settings.
We pre-train the model on the source with 60 percent of labeled nodes and adapt the model to the target. The adaptation involves three settings: (1) directly applying the pre-trained model without any fine-tuning (Without FT); (2) fine-tuning the last layer (classifier) of the model (FT Last Layer); (3) fine-tuning all parameters (Fully FT). We take split 10/10/80 to form train/valid/test sets on the target graph.
@@ 4.2 Node Classification
One-to-One Transfer.
The transfer learning results on Citation over three GNN backbones are presented in Table 2. Considering the limited number of parameters, we apply the FT Last Layer setting. Our GNN-SP outperforms baselines across all settings and is even better than the model trained from scratch (No Transfer), demonstrating the capability of overcoming negative transfer. Note that GNN-SP performs better than the advanced GNN-SP++ with extremely limited labels. It is may because (1) the dataset is small and sparse ($\sim$5k to $\sim$7k nodes and $\sim$20k to $\sim$30k edges) where RW sampler fails to model the real localized structures; and (2) the label sparsity exacerbates overfitting. The performance on Airport is presented in Table 3. Our method surpasses all baselines and achieves an average Rank of 1.5, which is notably higher than the No Transfer. Note that transferring knowledge to USA results in negative transfer for all baselines, likely due to its significantly larger size that potentially contains more patterns compared to the other two graphs. Specifically, USA contains 1,190 nodes and 28,388 edges, whereas Europe and Brazil contain only 399 nodes & 12,385 edges, and 131 nodes & 2,137 edges, respectively. The smaller graphs provide limited patterns and may introduce unexpected biases during pre-training. Additionally, we observe that GNN-SP++ performs better than GNN-SP, likely because these graphs are densely connected, leading to the issue that various nodes share identical $k$-hop subgraphs.
@@
One-to-Multi Transfer.
We use Twitch as the benchmark, which consists of six graphs with different sizes and data distributions. We pre-train the model on DE and fine-tune on other graphs (EN, ES, FR, PT, RU). We employ ROC-AUC as metric and adopt 2-layer GCN and SAGE as backbones. Figure 3 shows GNN-SP outperforms standard GNN with up to 8% improvements on ROC-AUC under FT Last Layer setting and achieves better performance than the model directly trained on the target over 10 out of 12 settings. The results validate the generalizability of GNN-SP to multiple graphs.
Transfer with Dynamic Shift.
In this scenario, we evaluate the model’s capability to handle temporal distribution shifts. We first adopt a dynamic financial network Elliptic with splitting 5/5/33 snapshots for train/valid/test. Figure 4 presents the results where the test snapshots are grouped into 9 folds. The performance gain of our GNN-SP is up to 10% and 24% over GAT and SAGE, respectively. Additionally, we use Arxiv as another temporal dataset, where nodes represent papers published from 2005 to 2020 and edges indicate citations. Based on the publication time, we collect five sub-graphs, represented as Time 1 (2005 - 2007), Time 2 (2008 - 2010), Time 3 (2011 - 2014), Time 4 (2015 - 2017), and Time 5 (2018 - 2020). We use the first four graphs as sources and the last one as the target. The results are presented in Table 4 where our GNN-SP++ achieves significant improvements over all baselines. We note that the temporal distribution shift is marginal when the source and target are temporally proximate, resulting in improved transfer learning performance. Additionally, the performance of GNN-SP++ is considerably superior to GNN-SP, further validating the efficacy of the RW sampler.